Exercise 14.8

The Pear Computer Company just developed a totally revolutionary new personal computer. Pear estimates that it will take competitors at least two years to produce equivalent products. The demand function for the computer is estimated to be

P=2,500−500QP=2,500−500Q

where QQ is millions of computers. The marginal (and average variable) cost of producing the computer is $900.

Assuming Pear acts as a monopolist in its market, the profit-maximizing price and output levels are

per computer and

million computers, respectively. The total contribution to profits and fixed costs at this output level is

million.

Pear Computer is considering an alternative pricing strategy of price skimming. It plans to set the following schedule of prices over the coming two years:

Complete the following table by calculating the contribution to profit and overhead for each of the 10 time periods and prices.

Time Period

Price

Quantity Sold

Total Contribution

($)

(Million)

($ Million)

1 2,400 0.2
2 2,200 0.2
3 2,000 0.2
4 1,800 0.2
5 1,700 0.2
6 1,600 0.2
7 1,500 0.2
8 1,400 0.2
9 1,300 0.2
10 1,200 0.2

Over the 10 periods, the total contribution to profits and fixed costs from price skimming is

million.
Answer #1

Given the demand function is P=2,500-5000. The marginal and average variable cost of producing the computer is $900. CalculatProfit maximizing point is where MR = MC MR = MC 2500 – 1000Q =900 1000Q =1,600 Q=1.6 Substitute Q in P: P=2,500-5009 = 2,500Below is the complete table: 2 A AVC 480 TC 900 900 9001 2200 4401 260 WN 0.21 Quantity Time Sold 1 Period Price ($) (MillionBelow table shows the formulas used for the calculation: A B C Time 1 Period 2 1 3 2 10.2 Quantity Sold Price ($) (Millions)